OpenStudy (anonymous):

Explain? Write an equation that describes the line passing through (4, -2) with a slope of 1/2? So I know that I have to put it into point slope form to get it into the slope intercept form, but how do you do that when you have fractions in it?

OpenStudy (anonymous):

Same as you would with any other numbers plug everything in, what do you have?

OpenStudy (anonymous):

So if you get y - ( - 2) = 1/2 (x - 4) how do you distribute the 1/2?

OpenStudy (anonymous):

like normal. 1/2 * x is (1/2)x what is 1/2 * -4?

OpenStudy (anonymous):

- 1?

OpenStudy (anonymous):

I meant - 2

OpenStudy (anonymous):

Yep so we have y+2=(1/2)x-2 now subtract 2 from both sides

OpenStudy (anonymous):

Okay so y + 2 -2 = (1/2)x - 2 - 2 = y = (1/2)x - 4?

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

You're welcome